Please help....
WORD PROBLEMS:Business and finance
Kevin earned $165 interest for 1 year on an investment of $1500. At the same rate, what amount of interest would be earned by an investment of $2500?

Answer by checkley71(8403)

I=P*R*T OR 165=1500*R*1 OR 1500R=165 OR R=165/1500 OR R=.11 OR AN INTEREST RATE OF 11% THUS I=2500*.11*1 OR I=275 FOR ONE YEAR.